# Intermediate Relation Size Bounds for Select-Project-Join Query Plans: Asymptotically Tight Characterizations

### v1

@@ Abstract.
We study the problem of statically optimizing select-project-join (SPJ) plans where unary key constraints are allowed. A natural measure of a plan, which we call the output degree and which has been studied previously, is the minimum degree of a polynomial bounding the plan’s output relation, as a function of the input database’s maximum relation size. This measure is, by definition, invariant under passing from a plan to another plan that is semantically equivalent to the first. In this article, we consider a plan measure which we call the intermediate degree; this measure is defined to be the minimum degree bounding the size of all intermediate relations computed during a plan’s execution — again, as a function of the input database’s maximum relation size. We present an algorithm that, given an SPJ plan $q$ and a set $\Sigma$ of unary keys, computes an SPJ plan $q^{\prime}$ that is semantically equivalent to $q$ (over databases satisfying $\Sigma$) and that has the minimum intermediate degree over all such semantically equivalent plans. For the types of plans considered, we thus obtain a complete and effective understanding of intermediate degree.
@@
Select-Project-Join Plans, Query Optimization, Tree Decompositions, Key Dependencies
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXX
1. Introduction
@@
Modern database systems capable of evaluating queries of the standard query language SQL typically have, as a key component, a query optimizer that synthesizes execution plans consisting of basic operations from a relational algebra; techniques for estimating and comparing the costs of evaluating such plans are considered central in the study of databases (Abiteboul et al., 1995, Chapter 6). In this article, we study a basic measure of execution plans: the intermediate size, by which we mean the maximum size over all intermediate relations computed during a query’s execution, relative to a database. We view this as a natural and fundamental measure: the intermediate size gives an obvious lower bound on the amount of time needed to execute a plan, for algorithms that explicitly materialize all intermediate relations; and, it characterizes the amount of space needed to store the results of all such intermediate relations.
@@
We study this size measure in an asymptotic setting, as a function of the maximum relation size $M_{\mathbf{D}}$ of a database $\mathbf{D}$. We say that a plan has intermediate degree $d$ when $d\geq 0$ and there exists a degree $d$ polynomial upper bounding the plan’s intermediate size, as a function of $M_{\mathbf{D}}$, but no lower degree polynomial provides such a bound. We study this measure for SPJ query plans, which allow select, project, and join as operations on relations. These plans are intensely studied throughout database theory. We permit unary key constraints to be expressed on database relations. In essence, a unary key constraint states that there is a single coordinate of a relation where, for each tuple in the relation, knowing the value at the coordinate determines the entire tuple.
@@
Our main theorem presents an algorithm that, given an SPJ plan $p$ with a set $\Sigma$ of unary key constraints, computes an SPJ plan $p^{\prime}$ that is semantically equivalent to $p$ (with respect to databases satisfying the constraints $\Sigma$), and has the lowest possible intermediate degree over all SPJ plans that are so semantically equivalent to $p$. This theorem thus yields, in the asymptotic setting considered, a full understanding of how to statically optimize the intermediate degree so that it is the minimum possible. We remark that (to our knowledge) this type of understanding was not previously known even for SPJ plans on databases without any constraints. We view the definition of the asymptotic notion of intermediate degree, and its supporting theory, as conceptual contributions of this work.
@@
Our result showing how to minimize the intermediate degree amounts to demonstrating how to compute, for any given plan, an asymptotically optimal algorithm in a restricted model of computation, namely, the class of SPJ plans. Our focus on polynomial degree is aligned with the notion of fixed-parameter tractability, whereby a problem is tractable essentially when there is a degree $d$ such that each so-called slice has an algorithm with running time bounded above by a degree $d$ polynomial; a slice is a set of all instances sharing the same parameter, which is a value associated with each instance.
@@
To prove our result, we make a number of technical contributions that we believe will be of significant utility in the future study of intermediate degree, which, again, we view as a highly natural measure of plan complexity. One tool that we present (in Section 4) is a theorem showing how to convert an SPJ query plan to a relational structure along with a certain form of tree decomposition of the structure; this allows us to apply structural measures based on tree decompositions to reason about the intermediate degree of plans. While the way that our main theorem’s algorithm computes the minimized plan $p^{\prime}$ from a given plan $p$ is, on a high level, quite natural—essentially, $p^{\prime}$ is based on the core of the chase of a relational structure naturally derived from $p$—the proof that $p^{\prime}$ is minimal in the desired sense is nontrivial, since we have to reason about arbitrary plans that are semantically equivalent to $p^{\prime}$. The core is, intuitively speaking, a minimized version of a relational structure; the chase is a well-known procedure that enforces constraints on a relational structure.
@@
Related work. Our work is inspired by and is closely related to the work of Atserias, Grohe, and Marx (Grohe and Marx, 2014; Atserias et al., 2013), as well as the follow-up work thereof by Gottlob, Lee, Valiant, and Valiant (Gottlob et al., 2012). A main result shown by (Atserias et al., 2013) yields a general upper bound—the AGM bound—on the output size of any pure join plan $q$, namely, a bound of $M_{\mathbf{D}}^{\rho^{*}(q)}$, where $\rho^{*}(q)$ is the so-called fractional edge cover number of the plan $q$. They also supply an argument that this upper bound is tight. Here, by a pure join plan, we mean a plan that consists only of joins and basic relations of a database.
@@
Let us define the output degree of a query to be the minimum degree of a polynomial bounding the query’s output relation, as a function of the maximum relation size $M_{\mathbf{D}}$ of a database $\mathbf{D}$. Note that while the output degree is clearly invariant under passing from a query to another query that is semantically equivalent to the first, the intermediate degree does not enjoy this invariance. The AGM bound along with the argument that it is tight establish $\rho^{*}(q)$ as the output degree of a pure join plan $q$. The article (Atserias et al., 2013) also exhibits, for each such pure join plan, a semantically equivalent join-project plan whose intermediate relations all have size less than (or equal to) the output size, implying a characterization of the minimum intermediate degree (over semantically equivalent plans) as the fractional edge cover number—for pure join plans. We view the article (Atserias et al., 2013) as begging the study of more general classes of plans beyond pure join plans: they prove explicitly the necessity of using projections in the plans witnessing that the intermediate relation sizes can be made to be at most the output size!
@@
Extending the AGM bound, the work (Gottlob et al., 2012) characterizes the output degree of any SPJ plan and considering databases where unary key constraints can be expressed; this is done by identifying a measure called the color number. Note that this measure generalizes the fractional edge cover number (as discussed in Section 4 of (Gottlob et al., 2012)). Our work builds on this measure and the accompanying understanding of unary key constraints. In particular, in the present work, we naturally extend the notion of color number to define a measure of tree decompositions. We remark that the results of (Gottlob et al., 2012) are phrased in terms of conjunctive queries, which are semantically equivalent to SPJ plans.
@@
An important and influential line of research emerging after the work of Atserias, Grohe, and Marx presented new join algorithms, so-called worst-case optimal join algorithms, with running times shown to be bounded above by the AGM bound—apart from a multiplicative overhead (see for example (Ngo et al., 2013, 2018; Ngo, 2018) and the discussion therein). Building on this work, further research presented algorithms, with running time analyses, for more general forms of queries, including queries with aggregates (Joglekar et al., 2016; Khamis et al., 2016). The focus of these further works are on presenting algorithms and upper bounding their running time in terms of structural measures, generally without considerations of semantic equivalence; the present article establishes a complementary result by showing, for each query plan $p$ (of the specified form), a tight lower bound on the running time of a restricted class of algorithms, namely, algorithms that can be described by plans that are semantically equivalent to the given plan $p$.
2. Preliminaries
@@
We use $\mathbb{Q}$ to denote the rational numbers, $\mathbb{Q}^{+}$ to denote the non-negative rational numbers, and $\mathbb{N}$ to denote the natural numbers, which we understand to include 00. For a natural number $k\in\mathbb{N}$, we use $[k]$ to denote the first $k$ positive natural numbers, that is, the set $\{i\in\mathbb{N}~{}|~{}1\leq i\leq k\}$. For a map $f:A\to B$ and some $S\subseteq A$, $f\upharpoonright S$ denotes the restriction of $f$ to $S$. For a tuple $\overline{a}=(a_{1},\ldots,a_{k})$ over a set $A$ and a map $f:A\to B$, by extension of notation, we write $f(\overline{a})$ for the entry-wise action of $f$ on the tuple, i.e., to denote $(f(a_{1}),\ldots,f(a_{k}))$; we further write $\{\overline{a}\}$ for the set $\{a_{1},\ldots,a_{k}\}$. When $A$ is a set, we use $A^{*}$ to denote the set of all finite-length tuples over $A$. When $\approx$ is a binary relation on a set $A$, we define $\approx^{*}$ as the reflexive-symmetric-transitive closure of $\approx$, that is, $\approx^{*}$ is the smallest equivalence relation on $A$ that contains $\approx$ as a subset.
2.1. Structures
@@
A signature $\sigma$ is a finite set of relation symbols with associated arity. We denote by $R/n$ that symbol $R$ has arity $n\in\mathbb{N}$, we may also write $\mathrm{ar}(R)$ for $n$. A structure $\mathbf{A}$ over a signature $\sigma$ consists of a finite set called the universe or domain of the structure, denoted $\mathsf{dom}(\mathbf{A})$, and, for each symbol $R\in\sigma$, a relation $R^{\mathbf{A}}$ over $\mathsf{dom}(\mathbf{A})$ whose arity is that of $R$. An isolated element of a structure $\mathbf{A}$ is a universe element $a\in\mathsf{dom}(\mathbf{A})$ such that $a$ does not appear in any tuple of any relation of $\mathbf{A}$. An open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$ is a pair consisting of a structure $\mathbf{A}$ over signature $\sigma$ with no isolated elements and a tuple $\overline{a}$ over the universe $\mathsf{dom}(\mathbf{A})$ of $\mathbf{A}$. For a structure $\mathbf{A}$ over signature $\sigma$, let $M_{\mathbf{A}}$ be the maximum size over all relations of $\mathbf{A}$, that is, $\max_{R\in\sigma}|R^{\mathbf{A}}|$.
@@
$\{h(\overline{a})~{}|~{}\mbox{$h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$}\}.$
@@
$\{h\upharpoonright S~{}|~{}\mbox{$h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$}\}.$
Example 2.1.
As a small example, let $\sigma$ be the signature containing a single relation $E$ of arity $2$, and let $\mathbf{A}_{0}$ be the structure with universe $\mathsf{dom}(\mathbf{A}_{0})=\{u,v_{1},v_{2}\}$ and with relation $E^{\mathbf{A}_{0}}=\{(u,v_{1}),(u,v_{2})\}$. Consider the open structure $(\mathbf{A}_{0},(u))$; when $\mathbf{D}$ is a directed graph, we have that $\mathsf{homs}(\mathbf{A}_{0},(u),\mathbf{D})$ is equal to the arity $1$ relation containing each vertex (of $\mathbf{D}$) having an outgoing edge. Next, consider the open structure $(\mathbf{A}_{0},(v_{1},v_{2}))$; when $\mathbf{D}$ is a directed graph, we have that $\mathsf{homs}(\mathbf{A}_{0},(v_{1},v_{2}),\mathbf{D})$ is equal to the arity $2$ relation containing each pair of vertices (of $\mathbf{D}$) that each receive an incoming edge from a common vertex. This relation is symmetric, and when $x$ is a vertex of $\mathbf{D}$, this relation contains the pair $(x,x)$ if and only if $x$ receives an incoming edge (from some vertex).
The next fact follows from the classic work of Chandra and Merlin (Chandra and Merlin, 1977).
Proposition 2.2 ((Chandra and Merlin, 1977)).
@@
For similar structures $\mathbf{A}$ and $\mathbf{B}$, we say that $\mathbf{A}$ is a substructure of $\mathbf{B}$ if $\mathsf{dom}(\mathbf{A})\subseteq\mathsf{dom}(\mathbf{B})$ and for each symbol $R\in\sigma$, it holds that $R^{\mathbf{A}}\subseteq R^{\mathbf{B}}$. When $\mathbf{B}$ is a structure and $S\subseteq\mathsf{dom}(\mathbf{B})$, the induced substructure $\mathbf{B}[S]$ is defined as the structure with universe $S$ and where, for each symbol $R$, it holds that $R^{\mathbf{B}[S]}=\{(b_{1},\ldots,b_{k})\in R^{\mathbf{B}}~{}|~{}\{b_{1},\ldots,b_{k}\}\subseteq S\}$. When $\mathbf{A}$ is a substructure of $\mathbf{B}$, a retraction from $\mathbf{B}$ to $\mathbf{A}$ is a homomorphism from $\mathbf{B}$ to $\mathbf{A}$ that acts as the identity on $\mathsf{dom}(\mathbf{A})$. When there exists a retraction from $\mathbf{B}$ to $\mathbf{A}$, we say that $\mathbf{B}$ retracts to $\mathbf{A}$. A core of a structure $\mathbf{B}$ is a substructure $\mathbf{C}$ of $\mathbf{B}$ such that $\mathbf{B}$ retracts to $\mathbf{C}$, but $\mathbf{C}$ does not retract to any proper substructure of $\mathbf{C}$, that is, any substructure of $\mathbf{C}$ whose universe is a proper subset of $\mathbf{C}$’s universe. A structure is a core if it is a core of itself.
Example 2.3.
Consider the structure $\mathbf{A}_{0}$ of Example 2.1. For each $i=1,2$, let $\mathbf{A}_{i}$ be the structure with universe $\mathsf{dom}(\mathbf{A}_{i})=\{u,v_{i}\}$ and with relation $E^{\mathbf{A}_{i}}=\{(u,v_{i})\}$. Clearly, each of $\mathbf{A}_{1},\mathbf{A}_{2}$ is a substructure of $\mathbf{A}_{0}$. In addition, we have that the map fixing $u$ and $v_{1}$ and sending $v_{2}$ to $v_{1}$ is a retraction from $\mathbf{A}_{0}$ to $\mathbf{A}_{1}$; dually, the map fixing $u$ and $v_{2}$ and sending $v_{1}$ to $v_{2}$ is a retraction from $\mathbf{A}_{0}$ to $\mathbf{A}_{2}$. Each of $\mathbf{A}_{1},\mathbf{A}_{2}$ is a core of $\mathbf{A}_{0}$ (and of itself): to argue this, we need to argue that (say) $\mathbf{A}_{1}$ has no retraction to a proper substructure $\mathbf{A}^{\prime}$ (of $\mathbf{A}_{1}$). (With respect to $\mathbf{A}_{0}$, the structures $\mathbf{A}_{1}$, $\mathbf{A}_{2}$ are symmetric to each other.) We argue this as follows: if we had such a proper substructure, it would have a size $1$ universe, and the retraction would map each element in $\mathsf{dom}(\mathbf{A}_{1})$ to the single element $x$ in that size $1$ universe. But then this proper substructure $\mathbf{A}^{\prime}$ would have $E^{\mathbf{A}^{\prime}}=\{(x,x)\}$, which could not be a subset of $E^{\mathbf{A}_{1}}$, contradicting that $\mathbf{A}^{\prime}$ is a substructure of $\mathbf{A}_{1}$.
@@
An open structure is not a structure per se. However, it will be useful to translate an open structure to a structure. We reserve a set $\{R_{k}\}_{k\in\mathbb{N}}$ of relation names for the purpose of translating open structures to structures, where each relation $R_{k}$ has arity $k$ and is distinct from the relations mentioned in any non-augmented signature. When $(\mathbf{A},(a_{1},\ldots,a_{k}))$ is an open structure over the signature $\sigma$, we define its augmented structure, denoted by $\mathsf{aug}(\mathbf{A},(a_{1},\ldots,a_{k}))$, as the structure $\mathbf{A}_{+}$ over signature $\sigma\cup\{R_{k}\}$ where $R^{\mathbf{A}_{+}}=R^{\mathbf{A}}$ for each $R\in\sigma$, and $R_{k}^{\mathbf{A}_{+}}=\{(a_{1},\ldots,a_{k})\}$. In particular, this means that if open structures $(\mathbf{A},\overline{a})$ and $(\mathbf{A}^{\prime},\overline{a^{\prime}})$ are similar and tuples $\overline{a}$ and $\overline{a^{\prime}}$ have the same arity, then also the augmented structures $\mathsf{aug}(\mathbf{A},\overline{a})$ and $\mathsf{aug}(\mathbf{A}^{\prime},\overline{a^{\prime}})$ are similar.
We associate each open structure with its augmented structure, so for example we regard an open structure $(\mathbf{C},\overline{c})$ as a core of an open structure $(\mathbf{B},\overline{b})$ when $\mathsf{aug}(\mathbf{C},\overline{c})$ is a core of $\mathsf{aug}(\mathbf{B},\overline{b})$. Note that a homomorphism from an open structure $(\mathbf{A},\overline{a})$ to an open structure $(\mathbf{B},\overline{b})$, as we defined it above, is a homomorphism from $\mathsf{aug}(\mathbf{A},\overline{a})$ to $\mathsf{aug}(\mathbf{B},\overline{b})$, and vice-versa.
2.2. Query plans
For completeness and clarity, we precisely define the types of select-project-join plans (SPJ plans) to be studied. For each $m\geq 0$, we define an $m$-suitable identification as an expression $j=k$ where $j,k\in[m]$.
For a signature $\sigma$, a SPJ plan over $\sigma$ is inductively defined as follows:
  • (basic) For each $R\in\sigma$, it holds that $R$ is a plan of arity $\mathrm{ar}(R)$.
  • (select) When $p$ is a plan of arity $m$ and $\theta$ is a set of $m$-suitable identifications, $\sigma_{\theta}(p)$ is a plan of arity $m$.
  • (project) When $p$ is a plan of arity $m$ and $j_{1},\ldots,j_{n}$ is a sequence of numbers from $[m]$, with $n\geq 0$, it holds that $\pi_{j_{1},\ldots,j_{n}}(p)$ is a plan of arity $n$.
  • (join) When $p_{1},\ldots,p_{\ell}$ are plans of arity $m_{1},\ldots,m_{\ell}$, respectively, with $\ell\geq 1$, and $\theta$ is a set of $(m_{1}+\cdots+m_{\ell})$-suitable identifications, $\Join_{\theta}(p_{1},\ldots,p_{\ell})$ is a plan of arity $(m_{1}+\cdots+m_{\ell})$.
Suppose that $Q$ is a relation of arity $m$. For any set $\theta$ of $m$-suitable identifications, we define
$\sigma_{\theta}(Q)=\{(d_{1},\ldots,d_{m})\in Q~{}|~{}\mbox{for each $(j=k)\in\theta$, $d_{j}=d_{k}$ holds}\}.$
For any sequence $j_{1},\ldots,j_{n}\in[m]$, we define
$\pi_{j_{1},\ldots,j_{n}}(Q)=\{(d_{j_{1}},\ldots,d_{j_{n}})~{}|~{}(d_{1},\ldots,d_{m})\in Q\}.$
Let $p$ be a SPJ plan over a signature $\sigma$, let $\mathbf{D}$ be a structure over $\sigma$, and let $m$ be the arity of $p$. We define the output of $p$ over $\mathbf{D}$, denoted by $\mathsf{out}(p,\mathbf{D})$, to be the arity $m$ relation over $D$ defined as follows. Note that, as is usual in this context, we understand that there is a unique tuple of arity 00, called the empty tuple.
  • When $p$ has the form $R$, we define $\mathsf{out}(p,\mathbf{D})=R^{\mathbf{D}}.$
  • When $p$ has the form $\sigma_{\theta}(p^{\prime})$, we define $\mathsf{out}(p,\mathbf{D})=\sigma_{\theta}(\mathsf{out}(p^{\prime},\mathbf{D})).$
  • When $p$ has the form $\pi_{j_{1},\ldots,j_{n}}(p^{\prime})$, we define $\mathsf{out}(p,\mathbf{D})=\pi_{j_{1},\ldots,j_{n}}(\mathsf{out}(p^{\prime},\mathbf{D})).$
  • When $p$ has the form $\Join_{\theta}(p_{1},\ldots,p_{\ell})$, we define $\mathsf{out}(p,\mathbf{D})=\sigma_{\theta}(\mathsf{out}(p_{1},\mathbf{D})\times\cdots\times\mathsf{out}(p_{\ell},\mathbf{D})).$

### v2

# Intermediate Relation Size Bounds for Select-Project-Join-Union Query Plans

@@ Abstract
We study the problem of statically optimizing select-project-join-union (SPJU) plans where unary key constraints are allowed. A natural measure of a plan, which we call the output degree and which has been studied previously, is the minimum degree of a polynomial bounding the plan’s output relation, as a function of the input database’s maximum relation size. This measure is, by definition, invariant under passing from a plan to another plan that is semantically equivalent to the first. In this article, we consider a plan measure which we call the intermediate degree; this measure is defined to be the minimum degree of a polynomial bounding the size of all intermediate relations computed during a plan’s execution — again, as a function of the input database’s maximum relation size. We present an algorithm that, given an SPJU plan $p$ and a set $\Sigma$ of unary keys, computes an SPJU plan $p^{\prime}$ that is semantically equivalent to $p$ (over databases satisfying $\Sigma$) and that has the minimum intermediate degree over all such semantically equivalent plans. For the types of plans considered, we thus obtain a complete and effective understanding of intermediate degree.
keywords:
@@
Select-Project-Join-Union Plans, Query Optimization, Tree Decompositions, Key Dependencies
1 Introduction
@@
Background and motivation. Modern database systems capable of evaluating queries of the standard query language SQL typically have, as a key component, a query optimizer that synthesizes execution plans consisting of basic operations from a relational algebra; techniques for estimating and comparing the costs of evaluating such plans are considered central in the study of databases [1, Chapter 6]. In this article, we study a basic measure of execution plans: the intermediate size, by which we mean the maximum size over all intermediate relations computed during a query’s execution, relative to a database. We view this as a natural and fundamental measure: the intermediate size gives an obvious lower bound on the amount of time needed to execute a plan, for algorithms that explicitly materialize all intermediate relations; and, it characterizes the amount of space needed to store the results of all such intermediate relations.
@@
We study this size measure in an asymptotic setting, as a function of the maximum relation size $M_{\mathbf{D}}$ of a database $\mathbf{D}$. We say that a plan has intermediate degree $d$ when $d\geq 0$ and there exists a degree $d$ polynomial upper bounding the plan’s intermediate size, as a function of $M_{\mathbf{D}}$, but no lower degree polynomial provides such a bound. We study this measure for SPJU query plans, which allow select, project, join and union as operations on relations. These plans are semantically equivalent to unions of conjunctive queries, and are intensely studied throughout database theory. We permit unary key constraints to be expressed on database relations. In essence, a unary key constraint states that there is a single coordinate of a relation where, for each tuple in the relation, knowing the value at the coordinate determines the entire tuple.
Having identified a relevant complexity measure of query plans—the intermediate degree—and a relevant class of query plans—the SPJU query plans, one of the most basic and natural questions one could ask, from the standpoint of query optimization, is this: is there an algorithm that receives a query plan as input, and reformulates it into a semantically equivalent query plan whose intermediate degree is minimal, over all such equivalent plans? That is, is there an algorithm that can statically analyze a given query plan, and reformulate it so as to optimally minimize the intermediate degree, without disrupting the plan’s semantics? A primary contribution of this article is a positive answer to this question.
@@
Contributions. Our main theorem presents a minimization algorithm that, given an SPJU plan $p$ with a set $\Sigma$ of unary key constraints, computes an SPJU plan $p^{\prime}$ that is semantically equivalent to $p$ (with respect to databases satisfying the constraints $\Sigma$), and has the lowest possible intermediate degree over all SPJU plans that are so semantically equivalent to $p$. This theorem thus yields, in the asymptotic setting considered, a full understanding of how to statically optimize the intermediate degree so that it is the minimum possible. We remark that (to our knowledge) this type of understanding was not previously known even for SPJ plans (SPJU plans that do not use the union operator) and on databases without any constraints. We view the definition of the notion of intermediate degree, and its supporting theory, as conceptual contributions of this work.
@@
Our result showing how to minimize the intermediate degree amounts to demonstrating how to compute, for any given plan, an asymptotically optimal procedure in a restricted model of computation, namely, the class of SPJU plans. Our focus on polynomial degree is aligned with the notion of fixed-parameter tractability, whereby a problem is tractable essentially when there is a degree $d$ such that each so-called slice has an algorithm with running time bounded above by a degree $d$ polynomial; a slice is a set of all instances sharing the same parameter, which is a value associated with each instance.
@@
To prove our result, we make a number of technical contributions that we believe will be of significant utility in the future study of intermediate degree, which we view as a highly natural measure of plan complexity. One tool that we present (in Subsection 4.1) is a theorem showing how to convert an SPJ query plan to a relational structure along with a certain form of tree decomposition of the structure; this allows us to apply structural measures based on tree decompositions to reason about the intermediate degree of plans. As a first step towards our main result, we show a restricted version of the main theorem to SPJ plans, that is, an algorithm that, given an SPJ plan $p$ and a set $\Sigma$ of unary keys, outputs a semantically equivalent SPJ plan $p^{\prime}$ (over all databases satisfying $\Sigma$) with the lowest possible intermediate degree. While the way that this algorithm computes the minimized plan $p^{\prime}$ from a given plan $p$ is, on a high level, quite natural — essentially, $p^{\prime}$ is based on the core of the chase of a relational structure naturally derived from $p$ — the proof that $p^{\prime}$ is minimal in the desired sense is non-trivial, since we have to reason about arbitrary plans that are semantically equivalent to $p^{\prime}$. Intuitively, the core is a minimized version of a relational structure and the chase is a well-known procedure that enforces constraints on a relational structure.
As defined, the intermediate degree is a measure of query plans. It follows from our definitions that any algorithm for evaluating a query plan $q$ that materializes all intermediate relations must require time at least $\Omega(M_{\mathbf{D}}^{d})$, where $d$ is the intermediate degree of $q$. That is, the polynomial time dependence on $M_{\mathbf{D}}$ must be of degree at least $d$. Having observed this lower bound, we are naturally led to the question of whether or not this bound is tight. We present results indicating that it is, which renders the intermediate degree as an algorithmically meaningful measure. When given a plan $p$, our minimization algorithm can always output a plan $p^{\prime}$ with the properties described above, and with the additional syntactic property that $p^{\prime}$ is well-behaved; as we show, well-behavedness of a plan $q$ implies that there is an evaluation algorithm for evaluating $q$ on a database $\mathbf{D}$ that always runs within time $M_{\mathbf{D}}^{d}$ times a multiplicative overhead, where $d$ is the intermediate degree of $q$. So, a well-behaved plan can be algorithmically evaluated in time close to $O(M_{\mathbf{D}}^{d})$, complementing the time lower bound of $\Omega(M_{\mathbf{D}}^{d})$ discussed above. We leave an analysis of the exact nature of the multiplicative overhead as a question for future work; we wish to emphasize here that the analysis of our evaluation algorithm and the resulting time bound are obtained in an elementary fashion.
@@
Related Work. Our work is inspired by and is closely related to the work of Atserias, Grohe, and Marx [12, 2], as well as the follow-up work thereof by Gottlob, Lee, Valiant, and Valiant [10]. A main result shown by [2] yields a general upper bound—the AGM bound—on the output size of any pure join plan $q$, namely, a bound of $M_{\mathbf{D}}^{\rho^{*}(q)}$, where $\rho^{*}(q)$ is the so-called fractional edge cover number of the plan $q$. They also supply an argument that this upper bound is tight. Here, by a pure join plan, we mean a plan that consists only of joins and basic relations of a database.
@@
Let us define the output degree of a query to be the minimum degree of a polynomial bounding the query’s output relation, as a function of the maximum relation size $M_{\mathbf{D}}$ of a database $\mathbf{D}$. Note that while the output degree is clearly invariant under passing from a query to another query that is semantically equivalent to the first, the intermediate degree does not enjoy this invariance. The AGM bound along with the argument that it is tight establish $\rho^{*}(q)$ as the output degree of a pure join plan $q$. The article [2] also exhibits, for each such pure join plan, a semantically equivalent join-project plan whose intermediate relations all have size less than (or equal to) the output size, implying a characterization of the minimum intermediate degree (over semantically equivalent plans) as the fractional edge cover number—for pure join plans. We view the article [2] as begging the study of more general classes of plans: they show explicitly the necessity of using projections in the plans witnessing that the intermediate relation sizes can be made to be at most the output size!
@@
Extending the AGM bound, the work [10] characterizes the output degree of any SPJ plan and considering databases where unary key constraints can be expressed; this is done by identifying a measure called the color number. Note that this measure generalizes the fractional edge cover number (as discussed in Section 4 of [10]). Our work builds on this measure and the accompanying understanding of unary key constraints. In particular, in the present work, we naturally extend the notion of color number to define a measure of tree decompositions. We remark that the results of [10] are phrased in terms of conjunctive queries, which are semantically equivalent to SPJ plans.
@@
An important and influential line of research emerging after the work of Atserias, Grohe, and Marx presented new join algorithms, so-called worst-case optimal join algorithms, with running times shown to be bounded above by the AGM bound—apart from a multiplicative overhead (see for example [19, 18, 17] and the discussion therein). Building on this work, further research presented algorithms, with running time analyses, for more general forms of queries, including queries with aggregates [13, 14]. The focus of these further works are on presenting algorithms and upper bounding their running time in terms of structural measures, generally without considering how to minimize queries up to semantic equivalence; the present article establishes a complementary result by showing, for each query plan $p$, a tight lower bound on the running time of a restricted class of procedures, namely, procedures that can be described by plans that are semantically equivalent to the given plan $p$.
Perspective. One measure of query plans that has been studied heavily is width, which by now has a mature underlying theory (refer for example to [15, 11, 6, 3, 20, 4, 7]); while this measure is often defined on logical formulas (and is sometimes formulated as the number of variables), for a query plan the width amounts to the maximum arity over all subplans. The present article is a contribution to a general research direction whereby we hope to identify further meaningful and informative measures of query plans, and study how to minimize various classes of queries with respect to introduced measures.
Outline. In Section 2, we present the necessary concepts used throughout the paper. Section 3 defines the central notion of intermediate degree, gives an intuition for the main theorem as well as discusses some algorithmic aspects of the plan output by the algorithm underlying the main theorem. In Section 4 we show how we can prove the main theorem restricted to SPJ plans before generalizing our result to SPJU plans in Section 5.
2 Preliminaries
@@
We use $\mathbb{Q}$ to denote the rational numbers, $\mathbb{Q}^{+}$ to denote the non-negative rational numbers, and $\mathbb{N}$ to denote the natural numbers, which we understand to include 0. For a natural number $k\in\mathbb{N}$, we use $[k]$ to denote the first $k$ positive natural numbers, that is, the set $\{i\in\mathbb{N}\penalty 10000\ |\penalty 10000\ 1\leq i\leq k\}$. For a map $f:A\to B$ and some $S\subseteq A$, $f\upharpoonright S$ denotes the restriction of $f$ to $S$. For a tuple $\overline{a}=(a_{1},\ldots,a_{k})$ over a set $A$ and a map $f:A\to B$, by extension of notation, we write $f(\overline{a})$ for the entry-wise action of $f$ on the tuple, i.e., to denote $(f(a_{1}),\ldots,f(a_{k}))$; we further write $\{\overline{a}\}$ for the set $\{a_{1},\ldots,a_{k}\}$. When $A$ is a set, we use $A^{*}$ to denote the set of all finite-length tuples over $A$. When $\approx$ is a binary relation on a set $A$, we define $\approx^{*}$ as the reflexive-symmetric-transitive closure of $\approx$, that is, $\approx^{*}$ is the smallest equivalence relation on $A$ that contains $\approx$ as a subset.
2.1 Structures
@@
A signature $\sigma$ is a finite set of relation symbols, each with an associated arity. We denote by $R/n$ that symbol $R$ has arity $n\in\mathbb{N}$; we may also write $\mathrm{ar}(R)$ for $n$. A structure $\mathbf{A}$ over a signature $\sigma$ consists of a finite set called the universe or domain of the structure, denoted $\mathsf{dom}(\mathbf{A})$, and, for each symbol $R\in\sigma$, a relation $R^{\mathbf{A}}$ over $\mathsf{dom}(\mathbf{A})$ whose arity is that of $R$. An isolated element of a structure $\mathbf{A}$ is a universe element $a\in\mathsf{dom}(\mathbf{A})$ such that $a$ does not appear in any tuple of any relation of $\mathbf{A}$. An open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$ is a pair consisting of a structure $\mathbf{A}$ over signature $\sigma$ with no isolated elements and a tuple $\overline{a}$ over the universe $\mathsf{dom}(\mathbf{A})$ of $\mathbf{A}$. For a structure $\mathbf{A}$ over signature $\sigma$, let $M_{\mathbf{A}}$ be the maximum size over all relations of $\mathbf{A}$, that is, $\max_{R\in\sigma}|R^{\mathbf{A}}|$.
@@
$\{h(\overline{a})\penalty 10000\ |\penalty 10000\ \mbox{$h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$}\}.$
@@
$\{h\upharpoonright S\penalty 10000\ |\penalty 10000\ \mbox{$h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$}\}.$
The next fact follows from the classic work of Chandra and Merlin [5].
Proposition 2.1 ([5]).
@@
For similar structures $\mathbf{A}$ and $\mathbf{B}$, we say that $\mathbf{A}$ is a substructure of $\mathbf{B}$ if $\mathsf{dom}(\mathbf{A})\subseteq\mathsf{dom}(\mathbf{B})$ and for each symbol $R\in\sigma$, it holds that $R^{\mathbf{A}}\subseteq R^{\mathbf{B}}$. When $\mathbf{B}$ is a structure and $S\subseteq\mathsf{dom}(\mathbf{B})$, the induced substructure $\mathbf{B}[S]$ is defined as the structure with universe $S$ and where, for each symbol $R$, it holds that $R^{\mathbf{B}[S]}=\{(b_{1},\ldots,b_{k})\in R^{\mathbf{B}}\penalty 10000\ |\penalty 10000\ \{b_{1},\ldots,b_{k}\}\subseteq S\}$. When $\mathbf{A}$ is a substructure of $\mathbf{B}$, a retraction from $\mathbf{B}$ to $\mathbf{A}$ is a homomorphism from $\mathbf{B}$ to $\mathbf{A}$ that acts as the identity on $\mathsf{dom}(\mathbf{A})$. When there exists a retraction from $\mathbf{B}$ to $\mathbf{A}$, we say that $\mathbf{B}$ retracts to $\mathbf{A}$. A core of a structure $\mathbf{B}$ is a substructure $\mathbf{C}$ of $\mathbf{B}$ such that $\mathbf{B}$ retracts to $\mathbf{C}$, but $\mathbf{C}$ does not retract to any proper substructure of $\mathbf{C}$, that is, any substructure of $\mathbf{C}$ whose universe is a proper subset of $\mathbf{C}$’s universe. A structure is a core if it is a core of itself. We provide an example illustrating the notions of (open) structure, the relation $\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$, substructure, and core for the interested reader in the appendix.
@@
[…]
2.2 Query plans